# When Relevance Meets Novelty: Dual-Stable Periodic Optimization for Exploratory Recommendation

### v1

@@ Abstract.
Traditional recommendation systems tend to trap users in strong feedback loops by excessively pushing content aligned with their historical preferences, thereby limiting exploration opportunities and causing content fatigue. Although large language models (LLMs) demonstrate potential with their diverse content generation capabilities, existing LLM-enhanced dual-model frameworks face two major limitations: first, they overlook long-term preferences driven by group identity, leading to biased interest modeling; second, they suffer from static optimization flaws, as a one-time alignment process fails to leverage incremental user data for closed-loop optimization. To address these challenges, we propose the Co-Evolutionary Alignment (CoEA) method. For interest modeling bias, we introduce Dual-Stable Interest Exploration (DSIE) module, jointly modeling long-term group identity and short-term individual interests through parallel processing of behavioral sequences. For static optimization limitations, we design a Periodic Collaborative Optimization (PCO) mechanism. This mechanism regularly conducts preference verification on incremental data using the Relevance LLM, then guides the Novelty LLM to perform fine-tuning based on the verification results, and subsequently feeds back the output of the incrementally fine-tuned Novelty LLM to the Relevance LLM for re-evaluation, thereby achieving a dynamic closed-loop optimization. Extensive online and offline experiments verify the effectiveness of the CoEA model in exploratory recommendation.
Exploratory Recommendation, Large Language Model
@@ 1. Introduction
  • Comprehensive Evaluation. Extensive online and offline experiments have demonstrated the effectiveness of our proposed framework for exploratory recommendation.
@@ 3.2. Causal Self-Attention
where the scaled dot product of query and key matrices is computed by dividing their product by $\sqrt{d_{k}}$. A causal mask $\bm{M}$ — with $\bm{M}_{ij}=0$ if $i\geq j$ and $\bm{M}_{ij}=-\infty$ if $i<j$ — is added to enforce autoregressive properties. Softmax is applied to this result, which is then multiplied by the value matrix to generate attention output $\bm{A}$. This mask prevents future information leakage, ensuring temporal consistency across $L$ layers. Each layer’s output is $\bm{H}^{(l)}=\text{LayerNorm}\left(\bm{A}+\bm{H}^{(l-1)}\right)$. For the final token at position $T$, its $l$ representation is $\bm{H}^{(l)}_{T}$; after $L$ layers, $\bm{H}^{(L)}_{T}$ encodes the entire sequence’s hierarchical context.
@@ 4.2. Dual-Stable Interest Exploration
Existing exploratory recommendation method (Wang et al., 2025a) mines novel items that users may be interested in by modeling their short-term behavior sequences. However, such methods fail to fully capture the long-term stable preferences implied by users’ group identities from short-term sequences. To address this issue, we propose the Dual-Stable Interest Exploration (DSIE) module
@@ 4.3.1. Novelty LLM Fine-tuning and Inference
[…]
@@ 4.3.3. Dynamic Closed-loop Optimization
[…]
@@ 5.4. Online A/B test (RQ3)
CoEA has been successfully deployed in Meituan App’s homepage recommendation system. As shown in Figure 5, the system processes $T-1$ historical behavior logs offline via Spark while capturing real-time behavior streams through Flink. Upon user request, the feature engine retrieves group CSIDs from offline User Table and item categories from Item Table, queries the Database for novel categories, and generates an additional retrieval path after filtering. Cold-start users are mapped to default group CSID with learning data updated daily. A 20-day A/B test demonstrates CoEA’s significant improvements: +1.203% GTV and +2.364% 7D-NIEP. This validates that CoEA effectively models the relationship between users’ long-term interests and novel categories, enabling it to discover latent interests while maintaining recommendation relevance, thereby achieving synergistic optimization of both business metrics and novelty.
@@ 5.5. Hyper-Parameter Studies (RQ4)
Impact of sequence filtering threshold. When $\tau$ is in the range of [5,9][5,9][ 5 , 9 ], the metrics C-H@1 and NCP@5 show fluctuations, where C-H@1 remains stable above 0.896 and NCP@5 remains stable above 0.152. When $\tau$ is in the range of [1,5][1,5][ 1 , 5 ], both C-H@1 and NCP@5 show an upward trend as $\tau$ increases, which indicates that noise interference will lead to a decrease in C-H@1 and C-N@5.
@@ 6. Conclusion and Future Work
Our proposed CoEA alleviates the Exploration-and-Exploitation dilemma in exploratory recommendation through two components: the first is DSIE, which can simultaneously capture the stable commonalities at the group level and the dynamic timeliness at the individual level; the second is PCO, which can dynamically optimize the generation of novel content based on real-time relevance signals. Experiments on industrial-scale datasets demonstrate the superior performance of CoEA: in offline evaluation, its recommendation quality and novelty discovery ability are significantly better than those of baseline models; the online deployment in Meituan App further verifies its practical value, successfully achieving a GTV increase of +1.203% and a 7D-NIEP growth of +2.364%. For future work, we will focus on optimizing the real-time performance of model updates for cold-start users.
@@ A.4.2. Category Normalized Discounted Cumulative Gain (C-N@K)
[…]
@@ A.4.3. Novel Category Proportion (NCP@K)
[…]
@@ A.4.4. Category Long-Tail Proportion (CLTP@K)
[…]
@@ A.4.5. Gross Transaction Value (GTV)
[…]
@@ A.6. Implementation Details
The sequence filtering click threshold is set to $\tau=5$; the CSA is configured with an embedding dimension of $d=128$, a layer depth of $L=4$, and 4 attention heads. For the Movielens-1M dataset, the configuration of RQ-VAE is as follows: the codebook size is 16, the number of layers of the residual quantizer is $l=4$, and the number of user groups $M_{N}$ is 50. For the MTRec dataset, the configuration of RQ-VAE is: the codebook size is 64, the number of layers of the residual quantizer is $l=4$, and the number of user groups $M_{N}$ is 4000. The number of group representative users is $k=8$, the DPO temperature coefficient is $\beta=0.1$, and the alignment threshold is $\tau_{\text{align}}=0.5$. The coefficient of $\alpha=0.4$ of KL divergence. The online model implements daily incremental fine-tuning and updates the offline database, while the offline model is configured differently based on datasets: 30 rounds for Movielens-1M and 1000 rounds for MTRec (each round consists of 5 batches with a batch size of 102410241024), with 1% of $\mathcal{D}_{\text{inc}}$ fed back to the Novelty LLM per epoch to synchronously optimize the Relevance LLM.

### v2

# Dual Collaborative LLMs via Continual Fine-Tuning for Serendipitous Recommendation

@@ Abstract.
Traditional recommendation systems tend to trap users in strong feedback loops by excessively pushing content aligned with their historical preferences, thereby limiting exploration opportunities and causing content fatigue. Although large language models (LLMs) demonstrate potential with their diverse content generation capabilities, existing LLM-enhanced dual-model frameworks face two major limitations: first, they overlook long-term preferences driven by group identity, leading to biased interest modeling; second, they suffer from static optimization flaws, as a one-time alignment process fails to leverage incremental user data for closed-loop optimization. To address these challenges, we propose the Co-Evolutionary Alignment (CoEA) method. For interest modeling bias, we introduce Dual-Stable Interest Exploration (DSIE) module, jointly modeling long-term group identity and short-term individual interests through parallel processing of behavioral sequences. For static optimization limitations, we design a Periodic Collaborative Optimization (PCO) mechanism. This mechanism regularly conducts preference verification on incremental data using the Relevance LLM, then guides the Novelty LLM to perform fine-tuning based on the verification results, and subsequently feeds back the output of the continually fine-tuned Novelty LLM to the Relevance LLM for re-evaluation, thereby achieving a dynamic closed-loop optimization. Extensive online and offline experiments verify the effectiveness of the CoEA model in serendipitous recommendation.
Serendipitous Recommendation, Large Language Model
@@ 1. Introduction
  • Comprehensive Evaluation. Extensive online and offline experiments have demonstrated the effectiveness of our proposed framework for serendipitous recommendation.
@@ 3.2. Causal Self-Attention
where the scaled dot product of query and key matrices is computed by dividing their product by $\sqrt{d_{k}}$. A causal mask $\bm{M}$ — with $\bm{M}_{ij}=0$ if $i\geq j$ and $\bm{M}_{ij}=$-$\infty$ if $i<j$ — is added to enforce autoregressive properties. Softmax is applied to this result, which is then multiplied by the value matrix to generate attention output $\bm{A}$. This mask prevents future information leakage, ensuring temporal consistency across $L$ layers. Each layer’s output is $\bm{H}^{(l)}=\text{LayerNorm}\left(\bm{A}+\bm{H}^{(l-1)}\right)$. For the final token at position $T$, its layer-$l$ representation is $\bm{H}^{(l)}_{T}$; after $L$ layers, $\bm{H}^{(L)}_{T}$ encodes the entire sequence’s hierarchical context.
@@ 4.2. Dual-Stable Interest Exploration
Existing serendipitous recommendation method (Wang et al., 2025a) mines novel items that users may be interested in by modeling their short-term behavior sequences. However, such methods fail to fully capture the long-term stable preferences implied by users’ group identities from short-term sequences. To address this issue, we propose the Dual-Stable Interest Exploration (DSIE) module
@@ 4.3.1. Novelty LLM Fine-tuning and Inference
where the fine-tuning prompt $\mathcal{P}^{(t-2:t-1)}_{\text{ft-nov}}$ combines the user group profile $\mathcal{D}_{s}$ with the (t2)(t-2)-th to (t1)(t-1)-th elements in the short-term category sequence $\mathcal{S}_{\text{short}}$, so as to predict the $t$-th short-term category $\mathcal{S}^{(t)}_{\text{short}}$. The objective function for SFT is:
@@ 4.3.3. Dynamic Closed-loop Optimization
where $\mathcal{P}$ represents the union of $\mathcal{P}_{\text{infer-nov}}$ and the incremental positive categories from the (k1)(k-1)-th cycle. $c_{\text{pos}}$ and $c_{\text{neg}}$ are the categories with the higher and lower relevance scores predicted by the Relevance LLM in the current $k$-th cycle, respectively. Based on the Direct Preference Optimization (DPO) method (Rafailov et al., 2023), incremental fine-tuning is performed on the Novelty LLM with reference to the fixed reference model $\pi_{\text{ref}}$ composed of the parameters of the Novelty LLM from the (k1)(k-1)-th cycle. Its optimization objective is:
@@ 5.4. Online A/B test (RQ3)
CoEA has been successfully deployed in Meituan App’s homepage recommendation system. As shown in Figure 5, the system processes T1T-1 historical behavior logs offline via Spark while capturing real-time behavior streams through Flink. Upon user request, the feature engine retrieves group CSIDs from offline User Table and item categories from Item Table, queries the Database for novel categories, and generates an additional retrieval path after filtering. Cold-start users are mapped to default group CSID with learning data updated daily. A 20-day A/B test demonstrates CoEA’s significant improvements: +1.203% GTV and +2.364% 7D-NIEP. This validates that CoEA effectively models the relationship between users’ long-term interests and novel categories, enabling it to discover latent interests while maintaining recommendation relevance, thereby achieving synergistic optimization of both business metrics and novelty.
@@ 5.5. Hyper-Parameter Studies (RQ4)
Impact of sequence filtering threshold. When $\tau$ is in the range of $[5,9]$, the metrics C-H@1 and NCP@5 show fluctuations, where C-H@1 remains stable above 0.896 and NCP@5 remains stable above 0.152. When $\tau$ is in the range of $[1,5]$, both C-H@1 and NCP@5 show an upward trend as $\tau$ increases, which indicates that noise interference will lead to a decrease in C-H@1 and C-N@5.
@@ 6. Conclusion and Future Work
Our proposed CoEA alleviates the Exploration-and-Exploitation dilemma in serendipitous recommendation through two components: the first is DSIE, which can simultaneously capture the stable commonalities at the group level and the dynamic timeliness at the individual level; the second is PCO, which can dynamically optimize the generation of novel content based on real-time relevance signals. Experiments on industrial-scale datasets demonstrate the superior performance of CoEA: in offline evaluation, its recommendation quality and novelty discovery ability are significantly better than those of baseline models; the online deployment in Meituan App further verifies its practical value, successfully achieving a GTV increase of +1.203% and a 7D-NIEP growth of +2.364%. For future work, we will focus on optimizing the real-time performance of model updates for cold-start users.
@@ A.4.2. Category Normalized Discounted Cumulative Gain (C-N@K)
(19) $\text{C-N@K}=\frac{1}{|\mathcal{U}|}\sum_{u\in\mathcal{U}}\frac{\text{DCG}_{u}}{\text{IDCG}_{u}}.$
@@ A.4.3. Novel Category Proportion (NCP@K)
(20) $\text{NCP@K}=\frac{1}{|\mathcal{U}|}\sum_{u\in\mathcal{U}}\frac{\text{new}_{u}}{|\mathcal{C}|}.$
@@ A.4.4. Category Long-Tail Proportion (CLTP@K)
(21) $\text{CLTP@K}=\frac{1}{|\mathcal{U}|}\sum_{u\in\mathcal{U}}\frac{\text{tail}_{u}}{|\mathcal{C}|}.$
@@ A.4.5. Gross Transaction Value (GTV)
(23) $\text{7D-NIEP}=\frac{1}{|\mathcal{U}|}\sum_{u\in\mathcal{U}}\frac{\text{new}_{u}}{\text{total}_{u}}.$
@@ A.6. Implementation Details
The sequence filtering click threshold is set to $\tau=5$; the CSA is configured with an embedding dimension of $d=128$, a layer depth of $L=4$, and 4 attention heads. For the Movielens-1M dataset, the configuration of RQ-VAE is as follows: the codebook size is 16, the number of layers of the residual quantizer is $l=4$, and the number of user groups $M_{N}$ is 50. For the MTRec dataset, the configuration of RQ-VAE is: the codebook size is 64, the number of layers of the residual quantizer is $l=4$, and the number of user groups $M_{N}$ is 4000. The number of group representative users is $k=8$, the DPO temperature coefficient is $\beta=0.1$, and the alignment threshold is $\tau_{\text{align}}=0.5$. The coefficient of $\alpha=0.4$ of KL divergence. The online model implements daily incremental fine-tuning and updates the offline database, while the offline model is configured differently based on datasets: 30 rounds for Movielens-1M and 1000 rounds for MTRec (each round consists of 5 batches with a batch size of $1024$), with 1% of $\mathcal{D}_{\text{inc}}$ fed back to the Novelty LLM per epoch to synchronously optimize the Relevance LLM.